OpenStudy (anonymous):

one leg of a right triangle is 2 more than twice the other. The hypotenuse is 13 meters. what is the lengths of the two legs

OpenStudy (anonymous):

Picture a triangle like this: They gave you hypotenuse = 13 c = 13 One leg is 2 more than twice the other: a = 2+2x The "other" leg: b = x Use the Pythagorean theorem to solve for x with the a, b and c values. \[a ^{2} + b ^{2} = c ^{2}\]

OpenStudy (anonymous):

Ok that you!! Got it So would I set it up as 2+2x+x=13

OpenStudy (anonymous):

Set it up like this: \[(x+2x)^{2} + x^{2} = 13^{2}\]

OpenStudy (anonymous):

sorry i mean \[(2+2x)^{2} + x^{2} = 13^{2}\]

OpenStudy (anonymous):

Got it.. Ok thank you!!

OpenStudy (anonymous):

Sure :)